Michael Scott's Paper Company common stock dividend is expected to grow at a long-run rate of 3% per year. The dividend recently paid was $1.50 per share. Investors require a 13% return from MSPC's common stock. What is your estimate of MSPC's common stock price?

If analysts suddenly change their estimate of MSPC's dividends growth rate to 6% instead of 3% what will happen toMSPC's stock price? The dividend recently paid was $1.50 per share. Investors require a 13% return from MSPC's common stock.
Answer #1

1)

Stock price D1:(r-g) Here, = 1.5*(1+3%) Expected dividend (D1) $ Required return (r) Growth rate (g) 1.55 13% 3% Stock price

2)

Stock price D1:(r-g) Here, = 1.5*(1+3%) Expected dividend (D1) $ Required return (r) Growth rate (g) 1.55 13% 6% Stock price

Please rate thumbs up